0peoplePeople, Dept of Computing, U of Newcastle0lion

——-

Steve Riddle

I am a Research Associate in the Centre for Software Reliability, working on the COPERNICUS project ISAT , Integration of Safety Analysis Techniques for Process Control Systems. Prior to arriving at Newcastle I was an undergraduate and then research student at the University of Bath, Computing Group. During my time at Bath I completed my first degree in Computer Software Technology, got married, spent a year as a Research Officer in the School of Mechanical Engineering looking at modelling the dynamics of underwater towed cable-and-body systems and back in the Computing Group began my PhD entitled "Viewpoints in Practice". This concerns the use of partial specifications and formal refinement theory to aid the process of explaining complex systems. My research interests include techniques for safety analysis, requirements engineering and formal methods. I am also interested in natural language processing and the use of computers in genealogy.

Publications

• M. Ainsworth, S. Riddle and PJL Wallis. Formal Validation of Viewpoint Specifications Software Engineering Journal January 1996 pp 58-66

See also my informal home page, and nameplate for address details.

Steve Riddle <steve.riddle@ncl.ac.uk>

Last modified: Wed Jul 17 08:09:16 BST 1996